# Detailed nuclear structure calculations for coherent elastic neutrino-nucleus scattering

### v1

@@ D.1 Argon-40
D.2 Germanium-70
@@ D.6 Germanium-76
D.7 Iodine-127
@@
D.8 Cesium-133
@@
The explicit nuclear form factors for 133Cs are given below using the convention of [35]:
@@
and can be used to obtain the $\mathcal{F}_{N}^{Y}(q^{2})$ in Eq. (IV.1). Since no interference occurs between operators in CEvNS, we simplify the notation and write $F^{(N,N^{\prime})}_{X,X}(q^{2})\equiv F^{(N,N^{\prime})}_{X}(q^{2})$.
@@
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.015-0.00813y+0.000808y^{2}+0.00263y^{3}-(8.21*10^{(}-6))y^{4}-0.0000477y^{5}+0.0000242y^{6}+5.25\times 10^{-7}y^{7}+4.03\times 10^{-9}y^{8})$
@@
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.00497+0.0599y-0.158y^{2}+0.164y^{3}-0.0838y^{4}+0.0234y^{5}-0.00359y^{6}+0.00028y^{7}-8.51\times 10^{-6}y^{8}+2.73\times 10^{-9}y^{9}$
@@
The occupation numbers $n^{N}_{nlj}$ of each valence single-nucleon orbital are given in order of increasing orbit, with $N=\{p,n\}$ denoting protons and neutrons respectively, and $n$ begins at $1$. “Max Occ” refers to the maximum occupation number allowed for each orbital. GCN5082* and SN100PN* refer to GCN5082 Expanded and SN100PN Expanded, respectively.
SN100PN SN100PN* GCN5082 GCN5082*

### v2

@@ D.1 Argon-40
The maximum difference between the form factors for the $M$ response at the minimum $q$ value is $\sim 0.3\%$, and $\sim 11\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 24\%$ and $\sim 21\%$, respectively, where the largest differences occur between the SDPF-MU and EPQQM interactions.
D.2 Germanium isotopes
The $M$ response is consistent amongst all calculations, with a maximum difference of $\sim 0.2\%$ at the lower end of the COHERENT $q$ region-of-interest (ROI) for the even isotopes, and $\sim 3\%$ at the higher end for 70,72Ge. For 73Ge these take on the values of $\sim 0.3\%$, and $\sim 20\%$ at the higher end for 73,74,76Ge. For the $\Phi^{\prime\prime}$ channel, the maximum difference is observed between the GCN2850 and jj44b shell model calculations, which is $\sim 80\%$ for 70Ge at both ends of the $q$ spectrum, factor of $\sim 2.4$ for 72Ge, factor of $\sim 2.3$ for 73Ge, factor of $\sim 2.2$ for 74Ge, and $\sim 60\%$ for 76Ge. The difference in the SD $\Sigma^{\prime}$ channel for 73Ge is $\sim 30\%$ at both ends of the $q$ ROI.
@@ D.2 Germanium isotopes
D.3 Iodine-127
The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 30\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 55\%$ and $\sim 50\%$, respectively. For $\Sigma^{\prime}$, the factor differences are $\sim 44$ and $\sim 81$, due to the large difference between the Truncated SN100PN calculation and the remaining calculations. This is consistent with the integrated form factor (IFF) results in reference [41], which showed a more pronounced difference between the SD Truncated SN100PN IFF results and other calculations. This indicates that a more severe valence space truncation for a shell model interaction has a larger effect compared to a second more comparable truncation.
@@ D.3 Iodine-127
D.4 Cesium-133
The maximum difference between the form factors for the $M$ channel at the minimum $q$ value is $\sim 1\%$, and $\sim 20\%$ at the maximum $q$ value. For $\Phi^{\prime\prime}$, these values are $\sim 6\%$ and $\sim 7\%$, respectively. For $\Sigma^{\prime}$, these differences are $\sim 13\%$ and $\sim 18\%$.
@@ D.4 Cesium-133
The explicit nuclear form factors for 133Cs are given below using the convention of Fitzpatrick et al. [38]:
@@
and can be used to obtain the $\mathcal{F}_{N}^{\Sigma^{\prime}}(q^{2})$ responses in Eq. (III.1) through the relation:
$|\mathcal{F}^{\Sigma^{\prime}}_{N}|^{2}(q^{2})=\frac{1}{4}\frac{2J_{i}+1}{4\pi}F_{\Sigma^{\prime}}^{N,N}$ (42)
This can be used to show that, in the isospin basis, the spin structure functions are:
$S^{\mathcal{T}}_{00}$ $=$ $\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{00}(\mathbf{q}^{2})$
$S^{\mathcal{T}}_{11}$ $=$ $(1+\delta^{\prime}(\mathbf{q}^{2}))^{2}\left(\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{11}(\mathbf{q}^{2})\right)$
$S^{\mathcal{T}}_{01}$ $=$ $2(1+\delta^{\prime}(\mathbf{q}^{2}))\left(\frac{2J_{i}+1}{4\pi}F^{\Sigma^{\prime}}_{01}(q^{2})\right)$
While describing CEvNS only requires the operators $M,\Phi^{\prime\prime}$ and $\Sigma^{\prime}$, below we provide the form factors for the full list of operators in [38].
@@
$F_{\tilde{\Phi}^{\prime}}^{(p,p)}=e^{-2y}(0.015-0.00813y+0.000808y^{2}+0.00263y^{3}-8.21\times 10^{-6}y^{4}-0.0000477y^{5}+0.0000242y^{6}+5.25\times 10^{-7}y^{7}+4.03\times 10^{-9}y^{8})$
@@
$F_{\Sigma^{\prime}\Delta}^{(n,p)}=e^{-2y}(-0.00497+0.0599y-0.158y^{2}+0.164y^{3}-0.0838y^{4}+0.0234y^{5}-0.00359y^{6}+0.00028y^{7}-8.51\times 10^{-6}y^{8}+2.73\times 10^{-9}y^{9})$
@@
The occupation numbers $n^{N}_{nlj}$ of each valence single-nucleon orbital are given in order of increasing orbit, with $N=\{p,n\}$ denoting protons and neutrons respectively, and $n$ begins at $1$. “Max Occ” refers to the maximum occupation number allowed for each orbital. “#2” refers the “Expanded” truncation schemes.
SN100PN SN100PN(#2) GCN5082 GCN5082(#2)